OpenStudy (anonymous):

solve for p--------- 2m-3p=1

jimthompson5910 (jim_thompson5910):

2m-3p=1 -3p=1-2m p=(1-2m)/(-3) p=(2m-1)/(3) So the answer is \[p=\frac{2m-1}{3}\]

OpenStudy (anonymous):

Woow,, you smart whats your name?

OpenStudy (anonymous):

The first step is to get everything with a \(p\) in it onto one side of the equation. You can either take \(2m\) from both sides; or add \(3p\), and take 1 from both sides.